# Data Trade and Consumer Privacy

### v2

@@ A.7 Proof of Proposition 5
$\pi\left(l|L\right),\pi\left(l|H\right)\in\left[0,1\right]$ ($FE$)
@@
Hence, there are two cases following the above Lemma: $(1)$ $IR_{L}$ is binding, $IR_{H}$ is not binding; $(2)$ both $IR_{L}$ and $IR_{H}$ are binding. Now we show Proposition 1 by three steps: Step 1 shows that it’s not optimal that only $IR_{L}$ is binding by contradiction; Step 2 simplifies the intermediary’s problem with binding $IR_{L}$ and $IR_{H}$; Step 3 solves the problem.
@@
00 0 $\geq$ $V-(x_{H}-L)^{2}-(x_{L}-H)^{2}+(x_{H}-H)^{2}+(x_{L}-L)^{2}-\left[\pi\left(l|L\right)+\pi\left(l|H\right)\right]\Delta v$
@@
We note that $IC_{L}$ is binding. To see this by contradiction, suppose the constraints ($IC_{L}$) are not binding. We have $\pi(l|L)=0$ as $v_{L}-\mu_{0}v_{H}<0$. Then $\pi(l|H)=0$ and $x_{L}=\frac{L-\mu_{0}H}{1-\mu_{0}}$ and $x_{H}=H$. But this contradicts with $IR_{H}$ in (22) as
@@
Suppose $x_{H}>H$. Then $\pi(l|H)=1$ and $\pi(l|L)=0$, which induces a contradiction with $OB_{L}$, $\pi(l|H)\leq\pi(l|L)$. Hence, $x_{H}=H$. Then $\pi(l|L)=0$. From the binding $IC_{L}$, $02\delta(x_{L}-x_{H})=(\pi(l|H)-\pi(l|L))\Delta v<$ 0, which implies $\pi(l|H)<0$, a contradiction.
@@
Its first order derivative with $\pi(l|L)$ is $-2\mu_{0}\left(x_{H}-H\right)\frac{(1-C)\Delta v}{2\delta}+(v_{L}-\mu_{0}v_{H})<0$, so $\pi(l|L)=0$ and then $\pi(l|H)=0$ by the binding $OB_{L}$. Hence, $x_{L}=x_{H}=L$ following the first order condition for $x_{L}$, i.e., $\mu_{0}(x_{L}-x_{H})-x_{L}+L=0$ and the binding $IC_{L}$: $02\delta(x_{L}-x_{H})=$ 0. It is a contradiction.
@@
$-\left[\frac{\delta}{2}-\frac{\pi(l|L)\Delta v}{2\delta}\right]^{2}+\pi\left(l|L\right)v_{L}$
@@
$\begin{cases}V+(1-\mu_{0})v_{L}&\mbox{ if }\frac{\delta^{2}}{\Delta v}\geq 1\\ V-(1-\mu_{0})\left(\frac{\Delta v-\delta^{2}}{2\delta}\right)^{2}+(1-\mu_{0})v_{L}&\mbox{ if }\frac{\Delta v}{v_{L}+v_{H}}\leq\frac{\delta^{2}}{\Delta v}<1\\ V+(1-\mu_{0})\left(\frac{\delta}{\Delta v}\right)^{2}v_{L}v_{H}&\mbox{ if }\frac{\delta^{2}}{\Delta v}<\frac{\Delta v}{v_{L}+v_{H}}\end{cases}$
@@
A.8 Additional Results on General Service
Proposition 7.
Suppose the service space $\mathcal{X}=\mathbb{R}$ and $\mu_{0}<\frac{v_{L}}{v_{H}}$. In the optimal mechanism, there is no distortion in service provisions $x_{L}=L$, $x_{H}=H$, the service fees are $f_{L}=V$, $f_{H}=V+[\pi(l|H)-1]\Delta v$, and the intermediary’s data trade is
$\pi\left(l|L\right)=1,\text{ }\pi\left(l|H\right)\in\Big{[}\max\Big{\{}0,1-\frac{\delta^{2}}{\Delta v}\Big{\}},1\Big{]}$
and the data trade fee is
$T=\mu_{0}[1-\pi\left(l|H\right)]\Delta v.$
In this case, social welfare is
$V+\left(1-\mu_{0}\right)v_{L}+\mu_{0}v_{H}.$
This actually goes back to the binary analysis in Proposition 2. The following proves Proposition 7.
@@ A.8 Additional Results on General Service
When $\mu_{0}<\frac{v_{L}}{v_{H}}$, the obedience constraint $OB_{L}$ is redundant, and the type-H consumers’ outside option max $\left\{v_{H}-p_{0},0\right\}=v_{H}-v_{L}$. Hence, the intermediary’s problem is simplified into
$\max_{f,x,\pi,T}f_{H}\mu_{0}+f_{L}\left(1-\mu_{0}\right)+T$
@@
$R\left(\pi\right)-R_{0}\geq T$ ($IR_{S}$)
$\pi\left(l|L\right),\pi\left(l|H\right)\in\left[0,1\right]$ ($FE$)
Lemma 7.
@@
Together with $f_{L}=V-\left(x_{L}-L\right)^{2}$, we have
@@
where the inequality follows from $IR_{H}$. So
$2x_{L}\geq L+H$
@@
Then $IC_{H}$ is
$2\left(x_{L}-x_{H}\right)\leq\pi\left(l|H\right)-\pi\left(l|L\right)$
and (non-binding) $IR_{L}$ is
$2x_{H}-v_{L}-v_{H}<1-\pi\left(l|H\right)$
@@
subject to $OB_{H}$ and $IC_{H}$
$2\left(x_{L}-x_{H}\right)\leq\pi\left(l|H\right)-\pi\left(l|L\right)$
and $\pi\left(l|L\right)$, $\pi\left(l|H\right)\in\left[0,1\right]$. If $IC_{H}$ is not binding and ignore the constraint, the solution is

### v3

@@ A.7 Proof of 3
$\pi\left(l|L\right),\pi\left(l|H\right)\in\left[0,1\right]$ (FEFE)
@@
Hence, there are two cases following the above Lemma: $(1)$ $IR_{L}$ is binding, $IR_{H}$ is not binding; $(2)$ both $IR_{L}$ and $IR_{H}$ are binding. Now we show 1 by three steps: Step 1 shows that it’s not optimal that only $IR_{L}$ is binding by contradiction; Step 2 simplifies the intermediary’s problem with binding $IR_{L}$ and $IR_{H}$; Step 3 solves the problem.
@@
$0$ $\geq$ $V-(x_{H}-L)^{2}-(x_{L}-H)^{2}+(x_{H}-H)^{2}+(x_{L}-L)^{2}-\left[\pi\left(l|L\right)+\pi\left(l|H\right)\right]\Delta v$
@@
We note that $IC_{L}$ is binding. To see this by contradiction, suppose the constraints ($IC_{L}$) are not binding. We have $\pi(l|L)=0$ as $v_{L}-\mu_{0}v_{H}<0$. Then $\pi(l|H)=0$ and $x_{L}=\frac{L-\mu_{0}H}{1-\mu_{0}}$ and $x_{H}=H$. But this contradicts with $IR_{H}$ in (24) because
@@
Suppose $x_{H}>H$. Then $\pi(l|H)=1$ and $\pi(l|L)=0$, which induces a contradiction with $OB_{L}$, $\pi(l|H)\leq\pi(l|L)$. Hence, $x_{H}=H$. Then $\pi(l|L)=0$. From the binding $IC_{L}$, $2\delta(x_{L}-x_{H})=(\pi(l|H)-\pi(l|L))\Delta v<0$, which implies $\pi(l|H)<0$, a contradiction.
@@
Its first order derivative with $\pi(l|L)$ is $2\mu_{0}\left(x_{H}-H\right)\frac{(1-C)\Delta v}{2\delta}+(v_{L}-\mu_{0}v_{H})<0$, so $\pi(l|L)=0$ and then $\pi(l|H)=0$ by the binding $OB_{L}$. Hence, $x_{L}=x_{H}=L$ following the first order condition for $x_{L}$, i.e., $\mu_{0}(x_{L}-x_{H})-x_{L}+L=0$ and the binding $IC_{L}$: $2\delta(x_{L}-x_{H})=0$. It is a contradiction.
@@
$\left[\frac{\delta}{2}-\frac{\pi(l|L)\Delta v}{2\delta}\right]^{2}+\pi\left(l|L\right)v_{L}$
@@
$\begin{cases}V+(1-\mu_{0})v_{L}&\mbox{ if }\frac{\delta^{2}}{\Delta v}\geq 1\\ V+(1-\mu_{0})v_{L}-(1-\mu_{0})\left(\frac{\Delta v-\delta^{2}}{2\delta}\right)^{2}&\mbox{ if }\frac{\Delta v}{v_{L}+v_{H}}\leq\frac{\delta^{2}}{\Delta v}<1\\ V+(1-\mu_{0})\left(\frac{\delta}{\Delta v}\right)^{2}v_{L}v_{H}&\mbox{ if }\frac{\delta^{2}}{\Delta v}<\frac{\Delta v}{v_{L}+v_{H}}\end{cases}$
@@
A.8 Proof of 4
@@
For $\mu_{0}>\frac{v_{L}}{v_{H}}$, social welfare with data banning is $V+\mu_{0}v_{H}$, and social welfare without regulation is
  1. (i)
    if $\frac{\delta^{2}}{\Delta v}\geq 1$: $V+(1-\mu_{0})v_{L}+\mu_{0}v_{H}$
  2. (ii)
  3. (iii)
Welfare in each case is strictly greater than $V+\mu_{0}v_{H}$. This is obvious for the first and third cases. In the second case, $v_{L}-\left(\frac{\Delta v-\delta^{2}}{2\delta}\right)^{2}=\frac{-(\Delta v)^{2}-\delta^{4}+2\delta^{2}(v_{L}+v_{H})}{4\delta^{2}}>0$ when $\frac{\Delta v}{v_{L}+v_{H}}<\frac{\delta^{2}}{\Delta v}<1$. Hence, banning data trade decreases social welfare for $\mu_{0}>\frac{v_{L}}{v_{H}}$.
A.9 Proof of 5
Proof.
The following completes the proof by two parts:
Proof of Part 1. of 5
@@ A.9 Proof of 5
When $\mu_{0}>\frac{v_{L}}{v_{H}}$, the obedience constraint for the producer to charge a high price $OB_{H}$ (18) is redundant, we can ignore (18). The type-H consumers’ outside option $\max\left\{v_{H}-p_{0},0\right\}=0$. So the intermediary’s problem is simplified into
$\max_{f,\pi}\text{ }f_{H}\mu_{0}+f_{L}\left(1-\mu_{0}\right)+\beta[\left(1-\mu_{0}\right)\pi\left(l|L\right)v_{L}-\mu_{0}\pi\left(l|H\right)\left(v_{H}-v_{L}\right)]$
s.t. (11)(13)(14)(15)(16)(17)
which are
$V-\left(x_{L}-L\right)^{2}-f_{L}\geq 0$ ($IR_{L}$)
$\pi\left(l|H\right)\leq C\pi\left(l|L\right)$ ($OB_{L}$)
$\pi\left(l|L\right),\pi\left(l|H\right)\in\left[0,1\right]$ (FEFE)
Lemma A.6.
If $IR_{H}$ (17) is binding, then $IR_{L}$ (14) is also binding.
Proof.
Suppose $IC_{L}$ is binding: $V-(x_{L}-L)^{2}-f_{L}=V-(x_{H}-L)^{2}-f_{H}$. From the binding $IR_{H}$ and $IC_{L}$, we have
$V-(x_{L}-L)^{2}-f_{L}$
$=$ $V-(x_{H}-L)^{2}-V+(x_{H}-H)^{2}-\pi\left(l|H\right)\Delta v$
$=$ $(L-H)(2x_{H}-L-H)-\pi\left(l|H\right)\Delta v$
$<$ $0$
Hence, there are two cases following the above Lemma: $(1)$ $IR_{L}$ is binding, $IR_{H}$ is not binding; $(2)$ both $IR_{L}$ and $IR_{H}$ are binding. Now we show in three steps: Step 1 shows that it’s not optimal that only $IR_{L}$ is binding by contradiction; Step 2 simplifies the intermediary’s problem with binding $IR_{L}$ and $IR_{H}$; Step 3 solves the problem.
Step 1.
also recall the binding $IR_{L}$: $V-(x_{L}-L)^{2}-f_{L}=0$, we have
$f_{L}=V-(x_{L}-L)^{2}$ (27)
Plugging (27) into (26) to get
Using $f_{L}$ in (27) and $f_{H}$ in (28), $IC_{L}$ becomes
$0$ $\geq$ $V-(x_{H}-L)^{2}-(x_{L}-H)^{2}+(x_{H}-H)^{2}+(x_{L}-L)^{2}-\left[\pi\left(l|L\right)+\pi\left(l|H\right)\right]\Delta v$
and ignore the (non-binding) $IR_{H}$
$(x_{L}-L)^{2}-(x_{L}-H)^{2}+\pi(l|L)\Delta v\geq 0$ (29)
$\quad\mu_{0}\left[V-(x_{L}-L)^{2}+(x_{L}-H)^{2}-(x_{H}-H)^{2}+(\pi(l|H)-\pi(l|L))\Delta v\right]$ (30)
$+(1-\mu_{0})\left[V-(x_{L}-L)^{2}\right]$
$+\beta[(1-\mu_{0})\pi(l|L)v_{L}-\mu_{0}\pi(l|H)\Delta v]$
$=$ $V-(x_{L}-L)^{2}+\mu_{0}\left[(x_{L}-H)^{2}-(x_{H}-H)^{2}\right]$
$\quad+\pi(l|L)[-\mu_{0}\Delta v+\beta(1-\mu_{0})v_{L}]+\pi(l|H)(1-\beta)\mu_{0}\Delta v$
subject to $IC_{L}$ and $OB_{L}$. It is useful to note $(1-\beta)\mu_{0}\Delta v>0$ and $\mu_{0}\Delta v+\beta(1-\mu_{0})v_{L}=\beta v_{L}-\mu_{0}(\Delta v+\beta v_{L})<0$ given $\mu_{0}>\frac{v_{L}}{v_{H}}$.
We also note that $IC_{L}$ is binding. To see this by contradiction, suppose the constraints ($IC_{L}$) are not binding and ignore it. There are two cases: (1) $OB_{L}$ is not binding. We have $\pi(l|L)=0$ as $\mu_{0}\Delta v+\beta(1-\mu_{0})v_{L}<0$, and $\pi(l|H)=1$ which however violates $OB_{L}$. (2) $OB_{L}$ is binding: $\pi(l|H)=C\pi(l|L)$. The objective function becomes
$V-(x_{L}-L)^{2}+\mu_{0}\left[(x_{L}-H)^{2}-(x_{H}-H)^{2}\right]$
$\quad+\pi(l|L)[-\mu_{0}\Delta v+\beta(1-\mu_{0})v_{L}]+C\pi(l|L)(1-\beta)\mu_{0}\Delta v$
$=$ $V-(x_{L}-L)^{2}+\mu_{0}\left[(x_{L}-H)^{2}-(x_{H}-H)^{2}\right]$
$\quad+\pi(l|L)[-\mu_{0}\Delta v+\beta(1-\mu_{0})v_{L}+C(1-\beta)\mu_{0}\Delta v]$
Then $\pi(l|L)=0$ as $\mu_{0}\Delta v+\beta(1-\mu_{0})v_{L}+C(1-\beta)\mu_{0}\Delta v<0$, and $x_{L}=\frac{L-\mu_{0}H}{1-\mu_{0}}$ and $x_{H}=H$. But this contradicts with $IR_{H}$ in (29) as
$\left(\frac{L-\mu_{0}H}{1-\mu_{0}}-L\right)^{2}-\left(\frac{L-\mu_{0}H}{1-\mu_{0}}-H\right)^{2}<0,$
so $IC_{L}$ is binding.
Using the binding $IC_{L}$, i.e.,
we have
$2\delta(x_{L}-x_{H})=\left[\pi(l|H)-\pi(l|L)\right]\Delta v$
from which by substituting out $x_{H}$, the objective in (30) becomes
$V-(x_{L}-L)^{2}+\mu_{0}\left[(x_{L}-H)^{2}-\left(x_{L}+\frac{(\pi(l|L)-\pi(l|H))\Delta v}{2\delta}-H\right)^{2}\right]$
$\quad+\pi(l|L)[-\mu_{0}\Delta v+\beta(1-\mu_{0})v_{L}]+\pi(l|H)(1-\beta)\mu_{0}\Delta v.$
Case (1): $OB_{L}$ is not binding. The first order derivatives with respect to $\pi(l|H)$ and $\pi(l|L)$ are
$-2\mu_{0}\left(x_{L}+\frac{(\pi(l|L)-\pi(l|H))\Delta v}{2\delta}-H\right)\frac{-\Delta v}{2\delta}+(1-\beta)\mu_{0}\Delta v$ $>$ $0$
$-2\mu_{0}\left(x_{L}+\frac{(\pi(l|L)-\pi(l|H))\Delta v}{2\delta}-H\right)\frac{\Delta v}{2\delta}-\mu_{0}\Delta v+\beta(1-\mu_{0})v_{L}$ $<$ $0$
where the second inequality follows $x_{H}\geq H$ and $\mu_{0}\Delta v+\beta(1-\mu_{0})v_{L}<0$ under $\mu_{0}>v_{L}/v_{H}$. So $\pi(l|H)=1$ and $\pi(l|L)=0$, which contradicts with $OB_{L}$.
Case (2): $OB_{L}$ is binding, i.e., $\pi(l|H)=C\pi(l|L)$. Then the objective function becomes
$V-(x_{L}-L)^{2}+\mu_{0}\left[(x_{L}-H)^{2}-\left(x_{L}+\frac{(\pi(l|L)-C\pi(l|L))\Delta v}{2\delta}-H\right)^{2}\right]$
$\quad+\pi(l|L)[-\mu_{0}\Delta v+\beta(1-\mu_{0})v_{L}]+C\pi(l|L)(1-\beta)\mu_{0}\Delta v$
Its first order derivative with $\pi(l|L)$ is
$2\mu_{0}\left(x_{L}+\frac{(\pi(l|L)-\pi(l|H))\Delta v}{2\delta}-H\right)\frac{(1-C)\Delta v}{2\delta}-\mu_{0}\Delta v+\beta(1-\mu_{0})v_{L}+C(1-\beta)\mu_{0}\Delta v<0$
where the inequality follows that $2\mu_{0}\left(x_{H}-H\right)\frac{(1-C)\Delta v}{2\delta}\leq 0$ and $\mu_{0}\Delta v+\beta(1-\mu_{0})v_{L}+C(1-\beta)\mu_{0}\Delta v<0$. So $\pi(l|L)=0$ and then $\pi(l|H)=0$ by the binding $OB_{L}$. Hence, $x_{L}=x_{H}$ following the binding $IC_{L}$. It is a contradiction.
Step 2.
$f_{H}$ $=$ $V-\left(x_{H}-H\right)^{2}+\pi\left(l|H\right)\Delta v$
$f_{L}$ $=$ $V-\left(x_{L}-L\right)^{2}$
Plugging $f_{L}$ and $f_{H}$ into the objective, the intermediary chooses $x_{H}$, $x_{L}$, $\pi\left(l|L\right)$, $\pi\left(l|H\right)$ to maximize
$f_{H}\mu_{0}+f_{L}\left(1-\mu_{0}\right)+\beta[\left(1-\mu_{0}\right)\pi\left(l|L\right)v_{L}-\mu_{0}\pi\left(l|H\right)\left(v_{H}-v_{L}\right)]$
$=$ $V-\mu_{0}\left(x_{H}-H\right)^{2}-(1-\mu_{0})\left(x_{L}-L\right)^{2}+(1-\beta)\mu_{0}\pi\left(l|H\right)\Delta v$
$+\beta\left(1-\mu_{0}\right)\pi\left(l|L\right)v_{L}$
$\pi\left(l|H\right)$ $\leq C\pi\left(l|L\right)$ ($OB_{L}$)
Step 3.
Solving the simplified problem:
$V-\mu_{0}\left(x_{H}-H\right)^{2}-(1-\mu_{0})\left(x_{L}-L\right)^{2}+(1-\beta)\mu_{0}\pi\left(l|H\right)\Delta v$
$+\beta\left(1-\mu_{0}\right)\pi\left(l|L\right)v_{L}$
subject to $\pi\left(l|L\right),\pi\left(l|H\right)\in\left[0,1\right]$, and
$\pi\left(l|H\right)$ $\leq C\pi\left(l|L\right)$ ($OB_{L}$)
Note that $x_{H}=H$ and $IC_{L}$ is redundant.
If $IC_{H}$ is not binding, then $x_{L}=L$, $\pi(l|L)=1$ and $\pi(l|H)=C$. $IC_{H}$ requires $\Delta v\leq\delta^{2}$.
If $\Delta v>\delta^{2}$, then $IC_{H}$ is binding, i.e., $\left(x_{L}-L\right)^{2}-\left(x_{L}-H\right)^{2}=-\pi\left(l|L\right)\Delta v$ from which
$(2x_{L}-L-H)\delta=-\pi(l|L)\Delta v$
or
$x_{L}=\frac{L+H}{2}-\frac{\pi(l|L)\Delta v}{2\delta}$
$V-(1-\mu_{0})\left(\frac{\delta}{2}-\frac{\pi(l|L)\Delta v}{2\delta}\right)^{2}+(1-\beta)\mu_{0}\pi\left(l|H\right)\Delta v+\beta\left(1-\mu_{0}\right)\pi\left(l|L\right)v_{L}$
Suppose $OB_{L}$ is not binding, then $\pi(l|H)=1$, contradicting with $OB_{L}$. So $OB_{L}$ is binding, and the objective function is
$V-(1-\mu_{0})\left(\frac{\delta}{2}-\frac{\pi(l|L)\Delta v}{2\delta}\right)^{2}+(1-\beta)\mu_{0}C\pi\left(l|L\right)\Delta v+\beta\left(1-\mu_{0}\right)\pi\left(l|L\right)v_{L}$
If $\pi(l|L)$ is not binding, the first order derivative with respect to $\pi(l|L)$ is
$2(1-\mu_{0})\left(\frac{\delta}{2}-\frac{\pi(l|L)\Delta v}{2\delta}\right)\frac{\Delta v}{2\delta}+(1-\beta)\mu_{0}C\Delta v+\beta\left(1-\mu_{0}\right)v_{L}=0$
from which
$\pi(l|L)$ $=$ $\frac{\delta^{2}}{\Delta v}+\frac{2(1-\beta)\mu_{0}C}{1-\mu_{0}}\frac{\delta^{2}}{\Delta v}+\frac{2\beta v_{L}}{\Delta v}\frac{\delta^{2}}{\Delta v}$
$=$ $\frac{\delta^{2}}{\Delta v}\left[1+\frac{2(1-\beta)\mu_{0}C}{1-\mu_{0}}+\frac{2\beta v_{L}}{\Delta v}\right]$
$=$ $\frac{\delta^{2}}{\Delta v}\frac{(1-\mu_{0})\Delta v+2\left[(1-\beta)\mu_{0}C\Delta v+\beta(1-\mu_{0})v_{L}\right]}{(1-\mu_{0})\Delta v}$
$=$ $\frac{\delta^{2}}{\Delta v}\frac{v_{H}+v_{L}}{\Delta v}$
where the last equality follows $\mu_{0}C\Delta v=(1-\mu_{0})v_{L}$ and the nonbinding $\pi(l|L)$ requires $\frac{\delta^{2}}{\Delta v}\frac{v_{H}+v_{L}}{\Delta v}\leq 1$ or $\frac{\delta^{2}}{\Delta v}\leq\frac{\Delta v}{v_{H}+v_{L}}$. Hence, the solution to this problem is
  • if $\frac{\delta^{2}}{\Delta v}\leq\frac{\Delta v}{v_{H}+v_{L}}$, then $\pi(l|L)=\frac{\delta^{2}}{\Delta v}\frac{v_{H}+v_{L}}{\Delta v}$ and $x_{L}=\frac{L+H}{2}-\frac{\Delta v}{2\delta}\pi(l|L)=L-\frac{\delta v_{L}}{\Delta v}$
  • if $\frac{\Delta v}{v_{H}+v_{L}}<\frac{\delta^{2}}{\Delta v}<1$, then $\pi(l|L)=1$ and $x_{L}=\frac{L+H}{2}-\frac{\Delta v}{2\delta}=L-\frac{\Delta v-\delta^{2}}{2\delta}$
In sum, we have $x_{H}=H$ and
  • if $\frac{\delta^{2}}{\Delta v}\geq 1$, then $\pi(l|L)=1$, $\pi(l|H)=C$ and $x_{L}=L$.
  • if $\frac{\Delta v}{v_{H}+v_{L}}<\frac{\delta^{2}}{\Delta v}<1$, then $\pi(l|L)=1$, $\pi(l|H)=C$ and $x_{L}=L-\frac{\Delta v-\delta^{2}}{2\delta}$.
  • if $\frac{\delta^{2}}{\Delta v}\leq\frac{\Delta v}{v_{H}+v_{L}}$, then $\pi(l|L)=\frac{\delta^{2}}{\Delta v}\frac{v_{H}+v_{L}}{\Delta v}$, $\pi(l|H)=\frac{C\delta^{2}}{\Delta v}\frac{v_{H}+v_{L}}{\Delta v}$ and $x_{L}=L-\frac{\delta v_{L}}{\Delta v}$.
In the optimal mechanism, the consumers surplus is 0 and the producer surplus is $\mu_{0}v_{H}$. The intermediary’s total revenue is
$\begin{cases}V+(1-\beta)\mu_{0}C\Delta v+\beta(1-\mu_{0})v_{L}&\mbox{ if }\frac{\delta^{2}}{\Delta v}\geq 1\\ V+(1-\beta)\mu_{0}C\Delta v+\beta(1-\mu_{0})v_{L}-(1-\mu_{0})\left(\frac{\Delta v-\delta^{2}}{2\delta}\right)^{2}&\mbox{ if }\frac{\Delta v}{v_{H}+v_{L}}\leq\frac{\delta^{2}}{\Delta v}<1\\ V+(1-\beta)\mu_{0}\frac{C\delta^{2}(v_{H}+v_{L})}{\Delta v}+\beta(1-\mu_{0})\frac{(v_{H}+v_{L})v_{L}\delta^{2}}{(\Delta v)^{2}}-(1-\mu_{0})\left(\frac{\delta v_{L}}{\Delta v}\right)^{2}&\mbox{ if }\frac{\delta^{2}}{\Delta v}<\frac{\Delta v}{v_{H}+v_{L}}\end{cases}$
By substituting out $C$, the above can be rewritten as
$\begin{cases}V+(1-\mu_{0})v_{L}&\mbox{ if }\frac{\delta^{2}}{\Delta v}\geq 1\\ V+(1-\mu_{0})v_{L}-(1-\mu_{0})\left(\frac{\Delta v-\delta^{2}}{2\delta}\right)^{2}&\mbox{ if }\frac{\Delta v}{v_{H}+v_{L}}\leq\frac{\delta^{2}}{\Delta v}<1\\ V+(1-\mu_{0})\left(\frac{\delta}{\Delta v}\right)^{2}v_{L}v_{H}&\mbox{ if }\frac{\delta^{2}}{\Delta v}<\frac{\Delta v}{v_{H}+v_{L}}\end{cases}$
which is independent of the intermediary’s bargaining power parameter $\beta$ and is the same as the base model.
Recall in the base setup with $\beta=1$, we have the following:
  • if $\frac{\delta^{2}}{\Delta v}\geq 1$, then $\pi(l|L)=1$ and $x_{L}=L$
The intermediary’s total revenue is also
$\begin{cases}V+(1-\mu_{0})v_{L}&\mbox{ if }\frac{\delta^{2}}{\Delta v}\geq 1\\ V-(1-\mu_{0})\left(\frac{\Delta v-\delta^{2}}{2\delta}\right)^{2}+(1-\mu_{0})v_{L}&\mbox{ if }\frac{\Delta v}{v_{L}+v_{H}}\leq\frac{\delta^{2}}{\Delta v}<1\\ V+(1-\mu_{0})\left(\frac{\delta}{\Delta v}\right)^{2}v_{L}v_{H}&\mbox{ if }\frac{\delta^{2}}{\Delta v}<\frac{\Delta v}{v_{L}+v_{H}}\end{cases}.$
Proof of Part 2. of 5
When $\mu_{0}<v_{L}/v_{H}$, the obedience constraint $OB_{L}$ is redundant and the type-H consumers’ outside option max $\left\{v_{H}-p_{0},0\right\}=v_{H}-v_{L}$. Hence, the intermediary solves
$\max_{f,x,\pi,T}f_{H}\mu_{0}+f_{L}\left(1-\mu_{0}\right)+\beta[R\left(\pi\right)-R_{0}]$
$\quad=f_{H}\mu_{0}+f_{L}\left(1-\mu_{0}\right)+\beta\left[\left(1-\mu_{0}\right)\pi\left(l|L\right)v_{L}-\mu_{0}\pi\left(l|H\right)\left(v_{H}-v_{L}\right)+\mu_{0}v_{H}-v_{L}\right]$
@@
$\frac{v_{L}}{v_{H}}\left(1-\mu_{0}\right)\pi\left(l|L\right)-(1-\frac{v_{L}}{v_{H}})\mu_{0}\pi\left(l|H\right)\geq\frac{v_{L}}{v_{H}}-\mu_{0}$ ($OB_{H}$)
$\pi\left(l|L\right),\pi\left(l|H\right)\in\left[0,1\right]$ (FEFE)
Lemma A.7.
@@
Together with the binding $IR_{L}$: $f_{L}=V-\left(x_{L}-L\right)^{2}$, we have
@@
where the inequality follows from $IR_{H}$. So the above inequality yields
$2x_{L}\geq L+H+[1-\pi(l|L)]\frac{\Delta v}{\delta}\geq L+H>2L$
@@
We show that the case (1) of only $IR_{H}$ (19) being binding is not feasible by contradiction. Suppose only $IR_{H}$ (19) is binding: $f_{H}=V-\left(x_{H}-H\right)^{2}+\left[\pi\left(l|H\right)-1\right]\Delta v$. As $IR_{L}$ is not binding, $IC_{L}$ must be binding. Otherwise, increasing $f_{L}$ will not affect $IR_{L}$, $IC_{L}$, $IC_{H}$ and will increase intermediary’s payoff. So the binding $IR_{H}$ (19) and $IC_{L}$ give
@@
Then substituting out $f_{H}$ and $f_{L}$ in $IC_{H}$, it becomes
$2\left(x_{L}-x_{H}\right)\leq\left[\pi\left(l|H\right)-\pi\left(l|L\right)\right]\frac{\Delta v}{\delta}$
and ignore the (non-binding) $IR_{L}$
$2x_{H}-L-H\leq\left[1-\pi\left(l|H\right)\right]\frac{\Delta v}{\delta}$
@@
$\max_{x,\pi}V-\left(x_{H}-{H}\right)^{2}$ $+(\pi(l|H)-1)\Delta v+\left[\left(x_{H}-{L}\right)^{2}-\left(x_{L}-{L}\right)^{2}\right]\left(1-\mu_{0}\right)$
$+\beta\left[\left(1-\mu_{0}\right)\pi\left(l|L\right)v_{L}-\mu_{0}\pi\left(l|H\right)\left(v_{H}-v_{L}\right)+\mu_{0}v_{H}-v_{L}\right]$
$=V-\left(x_{H}-{H}\right)^{2}$ $+\pi(l|H)(1-\beta\mu_{0})\Delta v+\pi(l|L)\beta(1-\mu_{0})v_{L}$
$+\left[\left(x_{H}-{L}\right)^{2}-\left(x_{L}-{L}\right)^{2}\right]\left(1-\mu_{0}\right)-(1-\beta\mu_{0})v_{H}-(1-\beta)v_{L}$
subject to $OB_{H}$ and $IC_{H}$: $2\left(x_{L}-x_{H}\right)\leq[\pi\left(l|H\right)-\pi\left(l|L\right)]\frac{\Delta v}{\delta}$, and $\pi\left(l|L\right)$, $\pi\left(l|H\right)\in\left[0,1\right]$. There are two cases depending on whether $IC_{H}$ is binding or not. We show $IC_{H}$ is binding by contradiction: Suppose $IC_{H}$ is not binding and ignore the constraint, the solution is